OpenStudy (poetryprincess):

If h(x) = 6 – x, what is the value of h*h)(10)? A. –4 B. –2 C. C.10 D. D. 16

OpenStudy (anonymous):

\[hh(10)=h(h(10))\]

OpenStudy (anonymous):

\[=h(6-10)\]

OpenStudy (anonymous):

\[=h(-4)\]

OpenStudy (anonymous):

\[=6-(-4)\]

OpenStudy (anonymous):

\[=6+4\]

OpenStudy (anonymous):

\[=?\]solve it @PoetryPrincess

OpenStudy (poetryprincess):

So the answer would be C. 10 correct? Thank you SO much.

OpenStudy (anonymous):

yep ^_^ yw! :)